Deming & Sons manufactures four grades of lubricant, W-10, W-20, W-30, and W-40, from a joint process. Additional informationProduct Joint Costs Allocated W-10 W-20 W-30 W-40
Answer #1
Joint costs allocated = Joint costs X Units produced of product/Total units produced
Product Joint costs allocated
W-10 133644 =381840*60200/172000
W-20 95460 =381840*43000/172000
W-30 76368 =381840*34400/172000
W-40 76368 =381840*34400/172000